The lengths (in minutes) of a random selection of popular children’s animated films are

listed below. Estimate the true mean length of all children’s animated films with 95%

confidence.

90 84 83 91 75 88 78 96 78 79 77


1
Expert's answer
2022-06-03T12:38:31-0400

Sample mean 


"\bar{x}=\dfrac{1}{11}(90+84+ 83+ 91 +75+ 88"




"+78 +96+ 78 +79+ 77)=\dfrac{919}{11}=84.515"

Sample variance


"s^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n-1}=47.072727272727""s=\sqrt{47.072727272727}\approx6.861"


The critical value for "\alpha = 0.05" and "df = n-1 = 10" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.2281"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{X}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{X}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(84.515-2.2281\times\dfrac{6.861}{\sqrt{11}},"

"84.515+2.2281\times\dfrac{6.861}{\sqrt{11}})"

"=(79.906,89.124)"

Therefore, based on the data provided, the 95 confidence interval for the population mean is "79.906 < \mu < 89.124," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(79.906, 89.124)."